# A path-dependent PDE solver based on signature kernels

### v3

@@ Abstract.
We develop a provably convergent kernel-based solver for path-dependent PDEs (PPDEs). Our numerical scheme leverages signature kernels, a recently introduced class of kernels on path-space. Specifically, we solve an optimal recovery problem by approximating the solution of a PPDE with an element of minimal norm in the signature reproducing kernel Hilbert space constrained to satisfy the PPDE at a finite collection of collocation paths. In the linear case, we show that the optimisation has a unique closed-form solution expressed in terms of signature kernel evaluations at the collocation paths. We prove consistency of the proposed scheme, guaranteeing convergence to the PPDE solution as the number of collocation points increases. Finally, several numerical examples are presented, in particular in the context of option pricing under rough volatility. Our numerical scheme constitutes a valid alternative to the ubiquitous Monte Carlo methods.
@@ 1.1. Path-dependent PDEs
The Feynman-Kac formula asserts that solutions to backward Kolmogorov equations have a probabilistic representation as $u_{0}(t,x)=\mathbb{E}[\phi(Y_{T})\lvert Y_{t}=x]$, for $0\leq t\leq T,\,x\in\mathbb{R}^{d}$, where $Y$ is an underlying Markov process. If $Y$ models the evolution of a financial asset price, then this expectation corresponds to the price of a European-type derivative written on this asset. In its seminal paper, Dupire [21] introduces a first generation of path-dependent PDEs (PPDEs) which solution is represented as $u_{1}(t,\gamma)=\mathbb{E}[\phi(Y_{[0,T]})\lvert Y_{[0,t]}=\gamma]$ where $\phi$ acts on the whole past of $Y$ and $\gamma$ is a path. This expectation represents the price of a path-dependent derivative. Rough volatility models, and more generally stochastic Volterra processes, are not Markov processes hence neither formulation applies. The recent framework of [54] tackles precisely this class of processes, for which the path-dependence is inherent to the dynamics. We illustrate this with an example to highlight the differences with the aforementioned cases. Let us consider a one factor rough volatility model
@@
where $X$ represents the log-price, $B$ and $W$ are correlated Brownian motions and generate their natural filtration $(\mathcal{F}_{t})_{t\in[0,T]}$, the kernel $K$ is square integrable and typically taken to be $K(t,r)=(t-r)^{H-\frac{1}{2}}$ with $H\in(0,1/2)$. In this situation, $\widehat{W}$ is a Gaussian Volterra process, which is neither a semimartingale nor a Markov process, hence why classical tools are not suitable. For $s\geq t$, we make the orthogonal decomposition $\widehat{W}_{s}=\Theta^{t}_{s}+I^{t}_{s}$ where $\Theta^{t}_{s}=\int_{0}^{t}K(s,r)\,\,\mathrm{d}W_{r}$ is $\mathcal{F}_{t}$-measurable while $I^{t}_{s}=\int_{t}^{s}K(s,r)\,\,\mathrm{d}W_{r}$ is independent from $\mathcal{F}_{t}$. As a consequence one recovers the Markovian structure as follows:
@@
This exhibits the path-dependence of the option price and entails there is a measurable map $u_{2}$ such that $u_{2}(t,x,\gamma)=\mathbb{E}\left[\phi(X_{T})|X_{t}=x,\Theta^{t}=\gamma_{[t,T]}\right]$. Building on Viens and Zhang’s functional Itô formula, further works [10, 55] study the associated path-dependent PDE of second generation of the form
(1.2) $\left\{\begin{array}[]{rl}&\mathcal{L}u(t,x,\gamma)=0,\\ &u(T,x,\gamma)=\phi(x),\end{array}\right.$
@@
Apart from the realm of rough volatility, stochastic Volterra processes have appeared in the modeling of electricity prices and turbulence [3, 7, 19, 39], in principal-agent problems [1], climate modeling [22, 52] and in asymptotic studies motivated by physcial phenomena [26, 27, 35]. More broadly, fractional Brownian motions and related processes are present in a variety of situations where intertemporal correlation is observed, with the seminal paper [37].
While the PPDE representation finds other applications (e.g. weak error rates in [10]), this article is concerned with their numerical computation. The solution to a PPDE is a function of a path (and other finite dimensional inputs). Contrary to the wide literature on numerical schemes for high-dimensional PDEs, we are here interested in purely infinite dimensional inputs that we will treat as such instead of discretising them. Despite the central place of such PPDEs in mathematical finance, numerical methods pertaining to them are scarse, see Section 1.6 for an overview of the literature. Our motivation is then twofold: design a numerical solver for such PDEs and overcome the technical challenges of learning a function on paths.
@@ 1.2. Signature kernels
Kernel methods are a well-established class of algorithms that are the key components of several machine learning models such as support vector machines (SVMs). The central idea of kernel methods is to lift the (possibly unstructured) input data to a (possibly infinite dimensional) Hilbert space by means of a nonlinear feature map defined in terms of a kernel. The advantage of doing so lies in the fact that a non-linear optimisation task in the original space becomes linear once lifted to the feature space. Thus, the solution of the original optimisation problem can often be expressed entirely in terms of evaluations of the kernel on training points by means of celebrated representer theorems. The kernel can often be evaluated efficiently with no reference to the feature map, a property known as kernel trick. The selection of an effective kernel will usually be a task-dependent problem, and this challenge is particularly difficult when the data is sequential. Signature kernels [33, 46] are a class of universal kernels on sequential data which have received attention in recent years thanks to their efficiency in handling path-dependent problems [34, 49, 17, 47, 16, 30, 29, 38].
@@ 1.3. Kernels vs neural networks for PDEs
In recent years, PDE solvers based on neural networks and kernel methods have been introduced as alternatives to classical finite difference and finite elements schemes. The most famous examples of neural PDE solvers are the Deep Galerkin model (DGM) [51] and Physics informed neural networks (PINNs) [43]. These techniques essentially parameterise the solution of the PDE as a neural network which is then trained using the PDE and its boundary conditions as loss function. More recently, DGM and PINNs have been extended to path-dependent PDEs by [45, 50, 31]. PDE solvers based on kernel methods have been investigated in various works [18, 12, 13], but the development of kernel methods for path-dependent PDEs has remained open. PDE solvers based on kernel methods hold potential for considerable advantages over neural PDE solvers, both in terms of theoretical analysis and numerical implementation. In effect, the theoretical analysis of neural PDE solvers is often limited to density or universal approximation results aimed at showing the existence of a network of a requisite size achieving a certain error rate, without guaranteeing whether this network is computable in practice via stochastic gradient descent. On the contrary, as we will demonstrate in this work, kernel-based PDE solvers are provably convergent and amenable to rigorous numerical analysis based on a rich arsenal of functional analytic results from the kernel literature.
@@ 1.4. Monte Carlo methods for pricing under rough volatility
The notion of rough volatility has recently disrupted stochastic modelling in quantitative finance [6], leaving no one indifferent in the field. In this setting, the instantaneous volatility of assets is modelled as a fractional Brownian motion with small Hurst parameter as in (1.1) or as a stochastic Volterra process. This feature has been shown empirically to be consistent with historical market trajectories and to capture important stylised facts of equity options in a parsimonous way. However, this new paradigm comes with an increased computational cost compared to classical techniques. With the notable exception of the rough Heston model, the absence of Markovianity of volatility prevents any pricing tools other than Monte Carlo simulations. However, the new promises–for estimation and calibration–of these rough volatility models have encouraged deep and fast innovations in numerical methods for pricing, in particular the now standard Hybrid scheme [8]. The complexity of a Monte Carlo scheme scales roughly as $\mathcal{O}(h^{-2}Nd^{2})$, where $N$ is the number of sample paths, $h$ is the discretisation step and $d$ is the dimension of the process. This complexity can be further reduced to $\mathcal{O}((h\log(h))^{-1}Nd^{2})$ with hybrid schemes [8]. In this paper, we propose a simulation-free alternative to Monte Carlo pricing by constructing a kernel-based solver for the corresponding PPDE. The complexity of the resulting solver scales as $\mathcal{O}(h^{-2}\hat{N}d^{2})$ where this time $\hat{N}$ is the number of collocation paths. This cost can be further reduced to $\mathcal{O}(h^{-1}\hat{N}d^{2})$ if the operations are carried out on a GPU, see [46] for details. As our numerical experiments demonstrate, a transparent comparison of complexities between our approach and Monte Carlo for pricing under rough volatility can only be achieved by obtaining error rates for the convergence of the approximations, which we intend to explore as future work.
@@ 1.5. Contributions
In this paper we design a novel numerical solver for PPDEs based on signature kernel and accompanied by convergence guarantees. Our numerical scheme does not use any classical probabilistic representation or any other pre-existing solver; it constitutes a valid alternative to the ubiquitous Monte Carlo methods, in particular in the context of option pricing under rough volatility. The hurdles in the theoretical analysis are many-fold and their resolution contributes to the development of the theory of signature kernels. Challenges include the translation of the PPDE framework into one amenable for the use of signature kernels, the proof of existence and uniqueness of a global minimum in the optimisation problem as well as an analytic form of the solution by means of a representer theorem, and a consistency result based on the adaptation of convergence results from [12] to this infinite-dimensional setting via a compactness argument for functions on path spaces (Theorem 4.4 and Lemma 4.6). Finally, several numerical examples are presented in Section 5 and the code is open-sourced at https://github.com/crispitagorico/sigppde.
@@ 1.6. Related literature
Several deep learning and signature-based approaches have been proposed in recent years. In [45], the authors use a probabilistic approach based on the martingale representation theorem to design an objective function for solving time-discretised PPDEs via a parametric model combining RNNs and signatures. The martingale representation theorem is also central in [23], where the authors make use of continuous-time RNNs developed in [40]. In both these papers, the signature has to be truncated for the implementation. The authors of [50] extend the deep Galerkin framework proposed by [51] using an LSTM architecture at each time step. The only paper applicable to Volterra processes [31] develops a backward discretisation algorithm using neural networks, drawing inspiration from the deep learning methodology introduced in [28]. We note that none of the aforementioned papers present theoretical results on the proposed numerical schemes and they all rely on the training of neural networks to approximate the solution. Aloof from the learning paradigm, convergence guarantees of monotone schemes can be found in [44] for the viscosity solutions to first generation PPDEs. The present paper is, to our knowledge, the first to propose a numerical scheme for second generation PPDEs with convergence guarantees.

### v4

@@ Abstract.
We develop a kernel-based solver for path-dependent PDEs (PPDEs) along with a convergence theory. Our numerical scheme leverages signature kernels, a recently introduced class of kernels on path-space. Specifically, we solve an optimal recovery problem by approximating the solution of a PPDE with an element of minimal norm in the signature reproducing kernel Hilbert space constrained to satisfy the PPDE at a finite collection of collocation paths. In the linear case, we show that the optimisation has a unique closed-form solution expressed in terms of signature kernel evaluations at the collocation paths. Under strict assumptions, we prove consistency of the proposed scheme, guaranteeing convergence to the PPDE solution as the number of collocation points increases. Finally, several numerical examples are presented, in particular in the context of option pricing under rough volatility. Our numerical scheme constitutes a valid alternative to the ubiquitous Monte Carlo methods.
@@ 1.1. Path-dependent PDEs
The Feynman-Kac formula asserts that solutions to backward Kolmogorov equations have a probabilistic representation as $u_{0}(t,x)=\mathbb{E}[\phi(Y_{T})\lvert Y_{t}=x]$, for $0\leq t\leq T,\,x\in\mathbb{R}^{d}$, where $Y$ is an underlying Markov process. If $Y$ models the evolution of the price of a financial asset, then this expectation corresponds to the price of a European-type derivative written on this asset. In its seminal paper, Dupire [21] introduces a first generation of path-dependent PDEs (PPDEs) which solution is represented as $u_{1}(t,\gamma)=\mathbb{E}[\phi(Y_{[0,T]})\lvert Y_{[0,t]}=\gamma]$ where $\phi$ acts on the whole past of $Y$ and $\gamma$ is a path. This expectation represents the price of a path-dependent derivative. Rough volatility models, and more generally stochastic Volterra processes, are not Markov processes hence neither formulation applies. The recent framework of [56] tackles precisely this class of processes, for which the path-dependence is inherent to the dynamics. We illustrate this with an example to highlight the differences with the aforementioned cases. Let us consider a one factor rough volatility model
@@
where $X$ represents the log-price, $B$ and $W$ are correlated Brownian motions and generate their natural filtration $(\mathcal{F}_{t})_{t\in[0,T]}$, the kernel $K$ is square integrable and typically taken to be $K(t,r)=(t-r)^{H-\frac{1}{2}}$ with $H\in(0,1/2)$. In this situation $\widehat{W}$ is a Gaussian Volterra process, which is neither a semimartingale nor a Markov process, that is why classical tools are not suitable. For $s\geq t$, we make the orthogonal decomposition $\widehat{W}_{s}=\Theta^{t}_{s}+I^{t}_{s}$ where $\Theta^{t}_{s}=\int_{0}^{t}K(s,r)\,\,\mathrm{d}W_{r}$ is $\mathcal{F}_{t}$-measurable while $I^{t}_{s}=\int_{t}^{s}K(s,r)\,\,\mathrm{d}W_{r}$ is independent from $\mathcal{F}_{t}$. As a consequence, one recovers the Markovian structure as follows:
@@
This exhibits the path-dependence of the conditional expectation $\mathbb{E}[\phi(X_{T})\lvert\mathcal{F}_{t}]$ (corresponding to an option price with underlying asset $\exp(X)$) and entails there is a measurable map $u$ such that $u(t,x,\gamma)=\mathbb{E}\left[\phi(X_{T})|X_{t}=x,\Theta^{t}_{[t,T]}=\gamma_{[t,T]}\right]$. Building on Viens and Zhang’s functional Itô formula, further works [10, 58] study the associated path-dependent PDE of second generation of the form
(1.2) $\left\{\hskip-11.38092pt\begin{array}[]{rl}&\mathcal{L}u(t,x,\gamma)=0,\\ &u(T,x,\gamma)=\phi(x),\end{array}\right$.
@@
Apart from the realm of rough volatility, stochastic Volterra processes have appeared in the modeling of electricity prices and turbulence [3, 8, 19, 39], in principal-agent problems [1], climate modeling [22, 54] and in asymptotic studies motivated by physical phenomena [26, 27, 35]. More broadly, fractional Brownian motions and related processes are present in a variety of situations where intertemporal correlation is observed, starting with the seminal paper [37].
While the PPDE representation finds other applications (e.g. weak error rates in [10]), this article is concerned with their numerical computation. The solution to a PPDE is a function of a path (and other finite-dimensional inputs). Contrary to the wide literature on numerical schemes for high-dimensional PDEs, we are here interested in purely infinite-dimensional inputs that we will treat as such instead of discretising them. Despite the central place of such PPDEs in mathematical finance, numerical methods pertaining to them are scarse; we refer to Section 1.6 for an overview of the literature. Our motivation is then twofold: design a numerical solver for such PDEs and overcome the technical challenges of learning a function on paths.
@@ 1.2. Signature kernels
Kernel methods are a well-established class of algorithms that are the key components of several machine learning models such as support vector machines (SVMs). The central idea of kernel methods is to lift the (possibly unstructured) input data to a (possibly infinite-dimensional) Hilbert space by means of a nonlinear feature map defined in terms of a kernel. The advantage of doing so lies in the fact that a non-linear optimisation task in the original space becomes linear once lifted to the feature space. Thus, the solution of the original optimisation problem can often be expressed entirely in terms of evaluations of the kernel on training points by means of celebrated representer theorems. The kernel can often be evaluated efficiently with no reference to the feature map, a property known as kernel trick. The selection of an effective kernel will usually be a task-dependent problem, and this challenge is particularly difficult when the data is sequential. Signature kernels [33, 48] are a class of universal kernels on sequential data which have received attention in recent years thanks to their efficiency in handling path-dependent problems [34, 51, 17, 49, 16, 30, 29, 38].
@@ 1.3. Kernels vs neural networks for PDEs
In recent years, PDE solvers based on neural networks and kernel methods have been introduced as alternatives to classical finite difference and finite elements schemes. The most famous examples of neural PDE solvers are the Deep Galerkin method (DGM) [53] and Physics informed neural networks (PINNs) [43]. These techniques essentially parameterise the solution of the PDE as a neural network which is then trained using the PDE and its boundary conditions as loss function. More recently, DGM and PINNs have been extended to path-dependent PDEs by [47, 52, 31]. PDE solvers based on kernel methods have been investigated in various works [18, 44, 45, 12, 13], but the development of kernel methods for path-dependent PDEs has remained open. PDE solvers based on kernel methods have the potential for considerable advantages over neural PDE solvers, both in terms of theoretical analysis and numerical implementation. In effect, the theoretical analysis of neural PDE solvers is often limited to density or universal approximation results aimed at showing the existence of a network of a requisite size achieving a certain error rate, without guaranteeing whether this network is computable in practice via stochastic gradient descent. In contrast, as we will demonstrate in this work, kernel-based PDE solvers are provably convergent under certain conditions and amenable to rigorous numerical analysis based on a rich arsenal of functional analytic results from the kernel literature. In addition to their theoretical appeal, kernel PDE solvers also enjoy concrete numerical benefits over neural PDE solvers. For linear PDEs such as those treated in this paper, for example, training reduces to solving a linear system of equations, whose stable numerical solutions are readily obtained with standard linear algebra routines. This stands in contrast to neural PDE solvers, where training involves non-convex optimization, often unstable in practice, and highly sensitive to hyperparameter choices.
@@ 1.4. Monte Carlo methods for pricing under rough volatility
The notion of rough volatility has recently disrupted stochastic modelling in quantitative finance [6], leaving no one indifferent in the field. In this setting, the instantaneous volatility of assets is modelled as a fractional Brownian motion with small Hurst parameter as in (1.1) or as a stochastic Volterra process. This feature has been shown empirically to be consistent with historical market trajectories and to capture important stylised facts of equity options in a parsimonious way. However, this new paradigm comes with an increased computational cost compared to classical techniques. With the notable exception of the rough Heston model, the absence of Markovianity of volatility prevents any pricing tools other than Monte Carlo simulations. However, the new promises––for estimation and calibration––of these rough volatility models have encouraged deep and fast innovations in numerical methods for pricing, in particular the now standard Hybrid scheme [7]. The complexity of a Monte Carlo scheme scales roughly as $\mathcal{O}(h^{-2}Nd^{2})$, where $N$ is the number of sample paths, $h$ is the discretisation step and $d$ is the dimension of the process. This complexity can be further reduced to $\mathcal{O}((h\log(h))^{-1}Nd^{2})$ with hybrid schemes [7]. In this paper, we propose a simulation-free alternative to Monte Carlo pricing by constructing a kernel-based solver for the corresponding PPDE. The complexity of the resulting solver scales as $\mathcal{O}(h^{-2}\hat{N}d^{2})$ where this time $\hat{N}$ is the number of collocation paths. This cost can be further reduced to $\mathcal{O}(h^{-1}\hat{N}d^{2})$ if the operations are carried out on a GPU, see [48] for details. As our numerical experiments demonstrate, a transparent comparison of complexities between our approach and Monte Carlo for pricing under rough volatility can only be achieved by obtaining error rates for the convergence of the approximations, which we intend to explore as future work.
@@ 1.5. Contributions
In this paper we design a novel numerical solver for PPDEs based on signature kernel and accompanied by a convergence theory. Our numerical scheme does not use any classical probabilistic representation or any other pre-existing solver; it constitutes a valid alternative to the ubiquitous Monte Carlo methods, in particular in the context of option pricing under rough volatility. Our kernel-based solver does not require additional “training” costs: the estimation of the Gram matrix is performed once, offline, and can then be reused for multiple configurations, payoffs, and evaluation points $(t,x,\gamma)$. By contrast, while Monte Carlo methods also allow the reuse of sample paths across different payoffs, a new simulation is required for each evaluation point $(t,x,\gamma)$. Moreover, Monte Carlo is only applicable when the PPDE admits a probabilistic representation. In settings where such a representation is unavailable, Monte Carlo cannot be used, whereas our solver remains applicable.
The hurdles in the theoretical analysis are many-fold and their resolution contributes to the development of the theory of signature kernels. Challenges include the translation of the PPDE framework into one amenable for the use of signature kernels, the proof of existence and uniqueness of a global minimum in the optimisation problem as well as an analytic form of the solution by means of a representer theorem, and a consistency result based on the adaptation of convergence results from [12] to this infinite-dimensional setting via a compactness argument for functions on path spaces (Theorem 4.5 and Lemma 4.7). Finally, several numerical examples are presented in Section 5 and the code is open-sourced at https://github.com/crispitagorico/sigppde.
@@ 1.6. Related literature
Several deep learning and signature-based approaches have been proposed in recent years. In [47], the authors use a probabilistic approach based on the martingale representation theorem to design an objective function for solving time-discretised PPDEs via a parametric model combining RNNs and signatures. The martingale representation theorem is also central in [23], where the authors make use of continuous-time RNNs developed in [40]. In both these papers, the signature has to be truncated for the implementation. The authors of [52] extend the deep Galerkin framework proposed by [53] using an LSTM architecture at each time step. The only paper applicable to Volterra processes [31] develops a backward discretisation algorithm using neural networks, drawing inspiration from the deep learning methodology introduced in [28]. We note that none of the aforementioned papers present theoretical results on the proposed numerical schemes and they all rely on the training of neural networks to approximate the solution. Aloof from the learning paradigm, convergence guarantees of monotone schemes can be found in [46] for the viscosity solutions to first generation PPDEs. The present paper is, to our knowledge, the first to propose a numerical scheme for second generation PPDEs with a convergence theory.
@@ 2.1. Path spaces
Let us set $d,e\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$, a finite time horizon $T>0$ and an interval $\mathbb{T}:=[0,T]$. Throughout this paper we deal with continuous paths $\gamma:\mathbb{T}\to\mathbb{R}^{e}$ and their perturbed càdlàg version $\gamma+\varepsilon\eta\mathbbm{1}_{[t,T]}$ for which we need to introduce a number of spaces and norms. When the interval $\mathbb{T}$ and state space $\mathbb{R}^{e}$ are clear from the context we will often omit them. The space of continuous functions from $\mathbb{T}$ to $\mathbb{R}^{e}$, denoted $\mathcal{C}(\mathbb{T},\mathbb{R}^{e})$, is equipped with the supremum norm $\left\lVert\gamma\right\rVert_{0;\mathbb{T}}:=\sup_{t\in\mathbb{T}}\left\lvert\gamma_{t}\right\rvert$, unless stated otherwise. The notation $\left\lvert\cdot\right\rvert$ refers to the Euclidean norm in one or multiple dimensions. The space of càdlàg paths from $\mathbb{T}$ to $\mathbb{R}^{e}$ is denoted $D(\mathbb{T},\mathbb{R}^{e})$ while for all $t\in\mathbb{T}$ we will make use of the space of càdlàg paths which are continuous on $[t,T]$, that is
$\mathcal{D}_{t}:=\Big\{\gamma\in D(\mathbb{T},\mathbb{R}^{e}):\gamma\big\lvert_{[t,T]}\in\mathcal{C}([t,T],\mathbb{R}^{e})\Big\}.$
This space allows to consider paths of the form $\gamma+\varepsilon\eta\mathbbm{1}_{[t,T]}$ where $\gamma,\eta\in\mathcal{C}(\mathbb{T},\mathbb{R}^{e})=\mathcal{D}_{0}$ and $\varepsilon>0$ which in turn permits the definition of pathwise derivatives on this space. In parallel, for $p\geq 1$, we also leverage the $p$-variation seminorm which is defined for $0\leq t<t^{\prime}\leq T$ and a path $\gamma:[t,t^{\prime}]\to\mathbb{R}^{e}$ as
$\left\lvert\gamma\right\rvert_{p-var;[t,t^{\prime}]}:=\left(\sup_{(t_{i})_{i}\in\pi}\sum_{i}\left\lvert\gamma_{t_{i+1}}-\gamma_{t_{i}}\right\rvert^{p}\right)^{1/p},$
where $\pi$ is the set of all partitions of the form $t=t_{0}<t_{1}<\cdots<t_{n}=t^{\prime}$ for some $n\in\mathbb{N}$. The associated norm is $\left\lVert\gamma\right\rVert_{p-var;[t,t^{\prime}]}=\left\lvert\gamma_{t}\right\rvert+\left\lvert\gamma\right\rvert_{p-var;[t,t^{\prime}]}$. This gives rise to the spaces $\mathcal{C}^{p-var}([t,t^{\prime}],\mathbb{R}^{e})$ of continuous paths with one strictly monotone coordinate  [^1] This is a technical assumption needed to ensure injectivity of the signature and consequently the point separating property assumption needed to establish universality of the signature kernel by Stone-Weierstrass arguments. More generally one could work with equivalence classes obtained by quotienting $\mathcal{C}$ by the so-called tree-like equivalence relation $x\sim_{\tau}y\iff S(x)=S(y)$; with the assumption that the paths have one strictly monotone coordinate, each equivalence class collapses to a singleton, which is the path itself. The monotone coordinate is usually taken to be time. and finite $p$-variation norm. For each such path $\gamma$ we denote by $\gamma^{0}$ the strictly monotone coordinate and by $\check{\gamma}$ the $\mathbb{R}^{e-1}$-valued path without the strictly monotone coordinate.
In the case $p=1$, the space $\mathcal{C}^{1-var}([t,t^{\prime}],\mathbb{R}^{e})$ is called the bounded variation space and also denoted $\mathrm{BV}([t,t^{\prime}],\mathbb{R}^{e})$. Of particular interest to us are the spaces $\mathcal{C}^{p-var}_{t}:=\mathcal{C}^{p-var}([t,T],\mathbb{R}^{e})$ and $\mathrm{BV}_{t}:=\mathcal{C}^{1-var}_{t}$ for some $t\in\mathbb{T}$. While the path space $\mathcal{D}_{t}$ in the original theory of [56] allows for one jump, this would lead to severe complications for the application of signature kernels. We show in the next subsection how to adapt our framework to avoid the introduction of jumps. Furthermore, we want to equip the space $\mathrm{BV}_{0}$ with a topology that renders the signature map continuous (see Section 3 for the details). We could choose any $1\leq p$-variation topology for this task, however the necessity to characterise compact sets of $\mathrm{BV}_{0}$ later on forces us to consider $p>1$. This task involves the introduction of Hölder spaces $\mathcal{C}^{\alpha}([t,t^{\prime}],\mathbb{R}^{e})$, with $\alpha\in(0,1)$ endowed with the $\alpha$-Hölder norm
If the interval $t,t^{\prime}]$ is clear from the context, for instance if $\gamma\in\mathrm{BV}([t,T],\mathbb{R}^{e})$, we omit to write it as a subscript. Having in mind to characterise compact subsets of $\mathcal{C}^{p-var}$, we follow [20] and intersect it with a Hölder space: let $\mathcal{C}^{p-var,\alpha}([t,t^{\prime}],\mathbb{R}^{e}):=\mathcal{C}^{p-var}([t,t^{\prime}],\mathbb{R}^{e})\cap\mathcal{C}^{\alpha}([t,t^{\prime}],\mathbb{R}^{e})$ be the space of $\alpha$-Hölder continuous paths with finite $p$-variation. The corresponding norm is $\left\lVert\gamma\right\rVert_{p-var,\alpha;[t,t^{\prime}]}:=\left\lvert\gamma\right\rvert_{\alpha;[t,t^{\prime}]}+\left\lvert\gamma\right\rvert_{p-var;[t,t^{\prime}]}$. It is then proven in [20, Theorem A.8] that the embedding $\mathcal{C}^{p-var,\alpha}\hookrightarrow\mathcal{C}^{p^{\prime}-var,\alpha^{\prime}}$ is compact for all $1\leq p<p^{\prime}\leq\infty$ and $0\leq\alpha^{\prime}<\alpha<1$ such that $\alpha p<1$ and $\alpha^{\prime}p^{\prime}<1$. In particular, bounded subsets of $\mathcal{C}^{1-var,\alpha}$ are compact with respect to the $p$-variation topology for all $\alpha\in(0,1)$ and $p>1$, see also [25, Proposition 5.30].
(2.1) $\widetilde{\Omega}=\Big\{(t,x,\gamma)\in\mathbb{T}\times\mathbb{R}^{d}\times\mathcal{D}_{t}\Big\},\qquad\widetilde{\bm{d}}((t,x,\gamma),(\bar{t},\bar{x},\bar{\gamma})):=\left\lvert t-\bar{t}\right\rvert+\left\lvert x-\bar{x}\right\rvert+\left\lVert\gamma-\bar{\gamma}\right\rVert_{0;\mathbb{T}}.$
Note that $\gamma$ lives in $\mathcal{D}_{t}$ which depends directly on $t$. We will denote by $\Omega$ the state space with $\mathrm{BV}_{0}$ instead of $\mathcal{D}_{t}$ and such that paths are constant on $[0,t]$ (except for the strictly monotone coordinate), that is
(2.2) $\Omega:=\big\{(t,x,\gamma)\in\mathbb{T}\times\mathbb{R}^{d}\times\mathrm{BV}_{0}:\check{\gamma}_{s}=\check{\gamma}_{0}\;\text{for all }s\leq t\big\}.$
The analogous distance is then
(2.3) $\bm{d}((t,x,\gamma),(\bar{t},\bar{x},\bar{\gamma})):=\left\lvert t-\bar{t}\right\rvert+\left\lvert x-\bar{x}\right\rvert+\left\lVert\gamma-\bar{\gamma}\right\rVert_{p-var;\mathbb{T}}.$
2.2. Functions of continuous paths